# Chaos and spatial prethermalization in driven-dissipative bosonic chains

### v2

@@ IV.2 Phase OTOC
$\{f,g\}:=\sum_{\ell=1}^{L}\left(\frac{\partial f}{\partial n_{\ell}}\frac{\partial g}{\partial\phi_{\ell}}-\frac{\partial g}{\partial n_{\ell}}\frac{\partial f}{\partial\phi_{\ell}}\right)\,.$ (13)
Carrying out this replacement in Eq. (12), using the relation $\{\phi_{\ell}(t^{\prime}),n_{k}(t)\}=-\delta\phi_{\ell}(t^{\prime})/\delta\phi_{k}(t)$, one obtains the semiclassical version of $C_{k,\ell}(t,\tau)$ which reads
$D_{k,\ell}(t,\tau)$ $=\frac{1}{2}\left\langle\left|\frac{\delta{\rm e}^{{\rm i}\phi_{\ell}(t+\tau)}}{\delta\phi_{k}(t)}\right|^{2}\right\rangle\,,$ (14)
where delimited-⟨$\langle$..$\rangle$ … ⟩ denote the average over the quantum noise and ${\delta}/{\delta\phi_{k}(t)}$ implements an infinitesimal perturbation of the phase at site $k$ and time $t$. In practice, this is implemented by cloning the system in two replicas $a$ and $b$, applying an infinitesimal perturbation at time $t$ on the phase at site $k$ of replica $b$, subsequently evolving both replicas subject to the same quantum noise, and finally averaging over realizations of the quantum noise. Thus, the semiclassical phase OTOC can be cast as
$D_{k,\ell}(t,\tau)=\frac{1}{2}\left\langle\left|{\rm e}^{{\rm i}\phi_{\ell}^{(a)}(t+\tau)}-{\rm e}^{{\rm i}\phi_{\ell}^{(b)}(t+\tau)}\right|^{2}\right\rangle\,,$ (15)
@@
Whenever the dynamics reaches a steady state, $\hat{\rho}_{\rm ss}:=\lim\limits_{t\to\infty}\hat{\rho}(t)$, we can define the semiclassical steady-state phase OTOC as $D^{\rm ss}_{k,\ell}(\tau):=\lim\limits_{t\to\infty}D_{k,\ell}(t,\tau)$. As a sanity check to back up the use of $D^{\rm ss}_{k,\ell}(\tau)$ as a dynamical probe of quantum chaos in the NESS, we verify that it reproduces the ballistic spreading of information one expects of a tight-binding model. Figure 6 clearly exhibits a causal light-cone structure with a finite butterfly velocity $v=2J$.
@@ Acknowledgements.
We acknowledge enlightening discussions with Alberto Biella, Andreas Läuchli, Alberto Mercurio, Luca Gravina, Lorenzo Fioroni and Pasquale Scarlino. CA acknowledges the support from the French ANR “MoMA” project ANR-19-CE30-0020 and the support from the project 6004-1 of the Indo-French Centre for the Promotion of Advanced Research (IFCPAR). FF, FM, and VS acknowledge the support by the Swiss National Science Foundation through Projects No. 200020_185015, 200020_215172, and 20QU-1_215928, and by the EPFL Science Seed Fund 2021.
@@ 1. Numerical approaches
The overall process leads to a stochastic Schrödinger equation for the wave function |$\ket{\psi(t)}$. Expectation values of operators can be obtained by averaging over many independent quantum trajectories (see the discussion below).
@@
The mapping of Eq. (1) to Eq. (Truncated Wigner approximation) is exact. The TWA consists in expanding the Moyal product ${\rm e}^{\overset{\leftrightarrow}{\Lambda}/2}$ up to second order and neglecting higher-order quantum fluctuations. In Eq. (Truncated Wigner approximation), this amounts to discarding the third-order derivative terms that stem from the Kerr nonlinearity while keeping lower-order contributions. This semiclassical approximation is valid in the limit of weak Kerr nonlinearity [40]. Within the TWA, the PDE becomes a Fokker-Planck equation [38] and $W$ can be interpreted as a well-defined probability distribution of the phase space variables. Finally, the Fokker-Planck equation can be mapped to a Langevin equation [38], yielding the set of stochastic differential equations (II.1). Expectation values of operators can be mapped from integrals over the Wigner function (22) into statistical expectation values of the Weyl symbols averaged over many Wigner trajectories solutions of Eqs. (II.1),
@@
To construct the steady-state quantity $W^{\rm ss}_{\ell}(\alpha,\alpha^{*}):=\lim\limits_{t\to\infty}W_{\ell}(t;\alpha,\alpha^{*})$, we realize a histogram from the data of $5\times 10^{7}$ fields $\alpha_{\ell}(t)$. The fields have been obtained from the time evolution of $20$ Wigner trajectories after average quantities have reached steady-state values, sampling over a time window $\Delta\tau=2\times 10^{5}$. Those numbers are fixed for all the Wigner functions presented in the manuscript.
The numerical results have been obtained by numerically solving Eqs. (II.1) using the solver SOSRI of the package Stochastic Differential Equations that is available in Julia. The initial conditions $\alpha_{\ell}(0)$ for each trajectory are complex random numbers sampled from a zero-mean Gaussian distribution with variance $1/2$ for $\ell=1,...,L$.
@@
${\rm i}\frac{\partial}{\partial t}\langle\hat{a}_{\ell}\rangle=-\Delta\langle\hat{a}_{\ell}\rangle+U\langle\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}\hat{a}_{\ell}\rangle-J\left(\langle\hat{a}_{\ell+1}\rangle+\langle\hat{a}_{\ell-1}\rangle\right)\,,$ (33)
@@
The QCE approach allows exploring quantum correlations beyond pure mean-field ansätze, but the set of equations for large system sizes becomes numerically intractable as we increase $d$. In particular, in the case of Eqs. (Quantum cumulant expansion), the number of cumulant equations scales as $L^{d}$. In this work, we focus on $d=2$, that captures Gaussian correlations. Calculations have been performed with the help of symbolic calculators for QCE equations available in Julia [78].
@@
Figure 7: Benchmark of TWA against exact dynamics (ED) for a driven-dissipative Bose-Hubbard chain described in Eq. (2) with length $L=2$. (a) Comparison between steady-state photon number $n$ obtained with TWA (colored lines) and Monte Carlo quantum trajectories (black markers) obtained by time evolving Eqs. (II.1) and the stochastic Schrödinger equation up to $t=100$. Averages have been performed over the time interval $[50,100]$ and over $500$ trajectories both for TWA and ED. The initial condition for TWA was $\alpha_{1}(0)=\alpha_{2}(0)=0$, while the initial state for ED was |$\ket{\psi(0)}=\ket{0}\otimes\ket{0}$. (b) Same analysis as in (a) but considering as initial states the coherent states $\alpha_{1}(0)=\alpha_{1}$, $\alpha_{2}(0)=\alpha_{2}$ and |$\ket{\psi(0)}=\ket{\alpha_{1}}\otimes\ket{\alpha_{1}}$, where $\alpha_{1}=\langle\hat{a}_{1}\rangle$ and $\alpha_{2}=\langle\hat{a}_{2}\rangle$ are computed with the TWA in the steady state. (c) Same analysis as in (a) but for the photon number statistics $4\delta n_{\ell}$. (d) Same analysis as in (b) but for the photon number statistics $\delta n_{\ell}$. ED and TWA show an excellent agreement in the vacuum, chaotic and high-photon number phases. The other parameters are set as in Fig. 3.
@@
In Fig. 8 we present the comparison between the TWA and the QCE. Figs. 8 (a) and (b) show $n_{L}^{\rm ss}$ and $\delta n_{L}^{\rm ss}$ computed solving cumulant equations truncated at order two. To compute the steady-state photon number and its fluctuations at each point in the phase diagram, we let the QCE dynamics evolve to the steady state, and we subsequently averaged the dynamics over a time window equal to $\Delta\tau=10^{3}$. The black line indicates the boundaries of the chaotic region computed from TWA equations. Data show a good qualitative agreement between TWA and QCE. However, as it is already evident from a rapid comparison between Figs. 3 (b) and (c) and Fig. 8, the two methods do not show a quantitative agreement.
@@
Figure 9: Saturation value of the steady-state phase OTOC, $D_{1,\ell}^{\rm ss}(\tau\to\infty)$ computed at different representative sites $\ell$ across the chain, as a function of the drive amplitude $F$ and the chain’s length $L$: (a) $\ell=1$, (b) $\ell=L/2$ , and (c) $\ell=L$ [same data as in Fig. 3 (a)]. The other parameters are set as in Fig. 3.
In this section, we provide additional plots to support the claim that the presence of quantum chaos is spread over the whole chain, and not only at the right boundary $\ell=L$ which is used in the main manuscript as a representative site. Notably, we find clear signatures of chaos in the prethermal domain, which justifies the term prethermal chaos that we used. In Fig. 9, we plot the saturation value of the phase OTOC $D_{1,\ell}^{\rm ss}(\tau\to\infty)$ as a function of the drive amplitude $F$ and the chain’s length $L$, when evaluated at three representative sites across the chain: first $\ell=1$, middle $\ell=L/2$, and last site $\ell=L$ (that was already presented in the main text).
When $\ell=1$, the saturation value never reaches 1, signaling that the first site does not reach maximal decorrelation. This is in agreement with the OTOC’s dynamics studied in Fig. 2 (a) for a specific value of $F$ and $L$. The phase diagram for $\ell=L/2$ is instead essentially identical to the one realized for $\ell=L$. This analysis shows that the prethermal and thermal domains are equally chaotic.
@@
In Fig. 10, we compare the boundary-driven boundary-dissipative chain studied in this paper and the uniformly-driven uniformly-dissipative chain, studied, e.g., in Ref. [40], showing that quantum chaos is a feature peculiar of boundary drive and dissipation mechanisms. We focus on the last cavities of the two chains in the steady-state regime. In Fig. 10 (a), we show that the photon number $n_{L}^{\rm ss}$ exhibits a similar behavior with an abrupt jump around the same critical drive amplitude $F$. In Fig. 10 (b), we present the photon number statistics $\delta n_{L}^{\rm ss}$. In the case of the uniformly-driven uniformly-dissipative chain, large fluctuations are concentrated at the value of $F$ where the jump occurs, as expected. In contrast, in the case of boundary drive and dissipation, large fluctuations are spread over a much larger region. In Fig. 10 (c), we plot the saturation value of the steady-state semiclassical phase OTOC defined in Eq. (5), showing final evidence that only the boundary-driven, boundary-dissipative lattice hosts quantum chaos, while the dynamics in the uniformly-driven dissipative chain is regular.
@@
Modeling the local state of the chaotic Bose-Hubbard chain with a U(1)-symmetric impurity model
@@
We consider two single-site driven-dissipative impurity models that describe U(1)-symmetric quantum states exhibiting, across their parameter space, thermal features and population inversion. Population inversion is achieved when higher energy states are more populated than lower energy states, in contrast to a thermal population, and requires an external energy source [79, 67, 80]. Population inversion is an essential ingredient to obtain lasing states, where stimulated emission dominates over absorption [81, 82]. The Hamiltonian of both the models is taken as the local Hamiltonian of the Bose-Hubbard chain in Eq. (2), $\hat{H}_{\rm imp}=\omega_{0}\hat{a}^{\dagger}\hat{a}+U\hat{a}^{\dagger 2}\hat{a}^{2}/2$. The first impurity model, which we refer as the 2-photon decay model, is defined by the set of Lindblad dissipators described in the main text,
$\hat{L}_{\uparrow}=\sqrt{\gamma^{\uparrow}}\hat{a}^{\dagger},\qquad\hat{L}_{\downarrow}=\sqrt{\gamma^{\downarrow}}\hat{a},\qquad\hat{L}_{\phi}=\sqrt{\gamma^{\phi}}\hat{a}^{\dagger}\hat{a},\qquad\hat{L}_{\rm s}=\sqrt{\gamma^{\rm s}}\hat{a}^{2}\,.$ (38)

### v3

@@ IV.2 Semiclassical OTOC
$\{f,g\}:=\sum_{\ell=1}^{L}\left(\frac{\partial f}{\partial n_{\ell}}\frac{\partial g}{\partial\varphi_{\ell}}-\frac{\partial g}{\partial n_{\ell}}\frac{\partial f}{\partial\varphi_{\ell}}\right)\,.$ (13)
Carrying out this replacement in Eq. (12), using the relation $\{\varphi_{\ell}(t^{\prime}),n_{k}(t)\}=-\delta\varphi_{\ell}(t^{\prime})/\delta\varphi_{k}(t)$, one obtains the semiclassical version of $C_{k,\ell}(t,\tau)$ which reads
$D_{k,\ell}(t,\tau)$ $=\frac{1}{2}\left\langle\left|\frac{\delta{\rm e}^{{\rm i}\varphi_{\ell}(t+\tau)}}{\delta\varphi_{k}(t)}\right|^{2}\right\rangle\,,$ (14)
where delimited-⟨$\langle$..$\rangle$ … ⟩ denote the average over the quantum noise and ${\delta}/{\delta\varphi_{k}(t)}$ implements an infinitesimal perturbation of the phase at site $k$ and time $t$. In practice, this is implemented by cloning the system in two replicas $a$ and $b$, applying an infinitesimal perturbation at time $t$ on the phase at site $k$ of replica $b$, subsequently evolving both replicas subject to the same quantum noise, and finally averaging over realizations of the quantum noise. Thus, the semiclassical phase OTOC can be cast as
$D_{k,\ell}(t,\tau)=\frac{1}{2}\left\langle\left|{\rm e}^{{\rm i}\varphi_{\ell}^{(a)}(t+\tau)}-{\rm e}^{{\rm i}\varphi_{\ell}^{(b)}(t+\tau)}\right|^{2}\right\rangle\,,$ (15)
@@
Whenever the dynamics reaches a steady state, $\hat{\rho}:=\lim\limits_{t\to\infty}\hat{\rho}(t)$, we can define the semiclassical steady-state phase OTOC as $D_{k,\ell}(\tau):=\lim\limits_{t\to\infty}D_{k,\ell}(t,\tau)$. Throughout the paper, we focus on $D_{k,\ell}(\tau)$ only.
We now show that our semiclassical OTOC is able to capture the basic features of quantum information spreading in extended many-body systems [90], as well as the initial exponential growth in the presence of chaos. On general grounds, we have $D_{k,\ell}(\tau=0)=0$, while $D_{k,\ell}(\tau>0)$ is expected to increase with $\tau$ whenever trajectories in the two replicas start deviating, and, eventually, to saturate to a finite value $D_{k,\ell}(\tau\to\infty)$. Both the growth and the saturation regimes of $D_{k,\ell}(\tau)$ shed light on the chaotic versus regular nature of the dynamics in the NESS. In Fig. 8 (a) we show the space-time dynamics of $D_{1,\ell}(\tau)$ in the chaotic regime of the boundary-driven dissipative Bose-Hubbard chain Eq. (2), for a system’s size equal to $L=80$. We see how $D_{1,\ell}(\tau)$ exhibits a causal light-cone structure with a ballistic spreading of information characterized by a butterfly velocity $v=2J$ [91]. In Fig. 8 (b) and (c), we illustrate the growth of the steady-state semiclassical OTOC for two representative values of the drive strength $F$, one for the chaotic regime, one for the regular RNW regime, always for a chain with $L=80$. In panel (b), $D_{1,\ell}(\tau)$ shows a rapid exponential growth of the form $D_{1,\ell}(\tau)\sim\exp[\lambda(t-\ell/v)]$ where $\lambda$ is a Lyapunov rate, followed by a saturation regime where $D_{1,\ell}(\tau\to\infty)\simeq 1$, i.e., maximal decorrelation. Altogether, the semiclassical OTOC $D_{1,\,\ell}(\tau)$ captures both the Lyapunov growth and the saturation regimes expected of quantum chaotic dynamics. In panel (c), $D_{1,\ell}(\tau)$ instead displays early-time oscillations and the overall growth is slower than exponential. Eventually, the late-time saturation value $D_{1,\ell}(\tau)$ is significantly less than $1$. This is indicative of regular dynamics.
@@ Acknowledgements.
We are grateful to Andreas Läuchli for invaluable suggestions on the analysis of local thermal equilibrium. We also acknowledge enlightening discussions with Alberto Biella, Abhishek Dhar, Lorenzo Fioroni, Luca Gravina, Manas Kulkarni, Alberto Mercurio, and Pasquale Scarlino. CA acknowledges the support from the French ANR “MoMA” project ANR-19-CE30-0020 and the support from the project 6004-1 of the Indo-French Centre for the Promotion of Advanced Research (IFCPAR). FF, FM, and VS acknowledge the support by the Swiss National Science Foundation through Projects No. 200020_185015, 200020_215172, and 20QU-1_215928, and by the EPFL Science Seed Fund 2021.
@@ 1. Numerical approaches
The overall process leads to a stochastic Schrödinger equation for the wave function |$\ket{\psi(t)}$. Expectation values of operators can be obtained by averaging over many independent quantum trajectories (see the discussion below). The numerical results for exact dynamics have been obtained with the QuantumToolbox.jl package [97] available in Julia.
@@
The mapping of Eq. (1) to Eq. (Truncated Wigner approximation) is exact. The TWA consists in expanding the Moyal product ${\rm e}^{\overset{\leftrightarrow}{\Lambda}/2}$ up to second order and neglecting higher-order quantum fluctuations. In Eq. (Truncated Wigner approximation), this amounts to discarding the third-order derivative terms that stem from the Kerr nonlinearity while keeping lower-order contributions. This semiclassical approximation is valid in the limit of weak Kerr nonlinearity [66]. Within the TWA, the PDE becomes a Fokker-Planck equation [56] and $W$ can be interpreted as a well-defined probability distribution of the phase space variables. Finally, the Fokker-Planck equation can be mapped to a Langevin equation [56], yielding the set of stochastic differential equations
${\rm i}\frac{\partial\alpha_{1}}{\partial t}$ $=-(\Delta+{\rm i}\gamma/2)\,\alpha_{1}+U\,(|\alpha_{1}|^{2}-1)\alpha_{1}-J\alpha_{2}+F+\sqrt{\frac{\gamma}{2}}\xi_{1}(t)\,,$
${\rm i}\frac{\partial\alpha_{\ell}}{\partial t}$ $=-\Delta\,\alpha_{\ell}+U\,(|\alpha_{\ell}|^{2}-1)\alpha_{\ell}-J(\alpha_{\ell-1}+\alpha_{\ell+1})\,,\quad\ell=2,...,L-1$ (31)
${\rm i}\frac{\partial\alpha_{L}}{\partial t}$ $=-(\Delta+{\rm i}\gamma/2)\,\alpha_{L}+U\,(|\alpha_{L}|^{2}-1)\alpha_{L}-J\alpha_{L-1}+\sqrt{\frac{\gamma}{2}}\xi_{L}(t)\,,$
coinciding with Eqs. (II.1). Expectation values of operators can be mapped from integrals over the Wigner function (22) into statistical expectation values of the Weyl symbols averaged over many Wigner trajectories solutions of Eqs. (II.1),
@@
To construct the steady-state quantity $W_{\ell}(\alpha,\alpha^{*}):=\lim\limits_{t\to\infty}W_{\ell}(t;\alpha,\alpha^{*})$, we realize a histogram from the data of $5\times 10^{7}$ fields $\alpha_{\ell}(t)$. The fields have been obtained from the time evolution of $10^{2}$ Wigner trajectories after average quantities have reached steady-state values, sampling over a time window $\Delta\tau=10^{4}$. Those numbers are fixed for all the Wigner functions presented in the manuscript.
The numerical results have been obtained by numerically solving Eqs. (II.1) using the solver SOSRI of the package Stochastic Differential Equations that is available in Julia. The initial conditions $\alpha_{\ell}(0)$ for each trajectory are complex random numbers sampled from a zero-mean Gaussian distribution with variance $1/2$ for $\ell=1,...,L$. The variance $1/2$ guarantees that the initial conditions correspond to the normalized vacuum state |$\outerproduct{0}{0}$ |.
@@
${\rm i}\frac{\partial}{\partial t}\langle\hat{a}_{\ell}\rangle=-\Delta\langle\hat{a}_{\ell}\rangle+U\langle\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}\hat{a}_{\ell}\rangle-J\left(\langle\hat{a}_{\ell+1}\rangle+\langle\hat{a}_{\ell-1}\rangle\right)\,,\quad\ell=2,...,L-1$ (34)
@@
The QCE approach allows exploring quantum correlations beyond the classical limit, but the set of equations for large system sizes becomes numerically intractable as we increase $d$. In particular, in the case of Eqs. (Quantum cumulant expansion), the number of cumulant equations scales as $L^{d}$. In this work, we focus on $d=2$, that captures Gaussian correlations. Calculations have been performed with the help of symbolic calculators for QCE equations available in Julia [99].
Gross-Pitaevksii classical equations of motion
The classical limit of the Lindblad equation Eq. (1) is described in terms of the zero-noise limit of the Langevin equations (Truncated Wigner approximation). From the TWA, the classical limit is obtained by reducing the Moyal product ${\rm e}^{\overset{\leftrightarrow}{\Lambda}/2}$ to the standard (commutative) product. The equations reduce to the classical Gross-Pitaevskii equations of motion reading [49]
${\rm i}\frac{\partial\alpha_{1}}{\partial t}=-\left(\Delta+{\rm i}\gamma/2\right)\alpha_{1}+U|\alpha_{1}|^{2}\alpha_{1}\ -J\alpha_{2}+F\,,$
${\rm i}\frac{\partial\alpha_{\ell}}{\partial t}=-\Delta\alpha_{\ell}+U|\alpha_{\ell}|^{2}\alpha_{\ell}-J\left(\alpha_{\ell+1}+\alpha_{\ell-1}\right)\,,\quad\ell=2,...,L-1$ (35)
${\rm i}\frac{\partial\alpha_{L}}{\partial t}=-\left(\Delta+{\rm i}\gamma/2\right)\alpha_{L}+U|\alpha_{L}|^{2}\alpha_{L}-J\alpha_{L-1}\,.$
They describe the time evolution of the field’s coherence $\alpha_{\ell}(t)$, starting initial conditions $\alpha_{\ell}(0)$ in the complex plane. The numerical solution of Eqs. (Gross-Pitaevksii classical equations of motion) is obtained using the package Differential Equations that is available in Julia.
@@
Figure 9: Benchmark of TWA against exact dynamics (ED) for a driven-dissipative Bose-Hubbard chain described in Eq. (2) with length $L=2$. (a) Comparison between steady-state photon number $n$ obtained with TWA (colored lines) and Monte Carlo quantum trajectories (black markers) obtained by time evolving Eqs. (II.1) and the stochastic Schrödinger equation up to $t=100$. Averages have been performed over the time interval $[50,100]$ and over $500$ trajectories both for TWA and ED. The initial condition for TWA was $\alpha_{1}(0)=\alpha_{2}(0)=0$, while the initial state for ED was |$\ket{\psi(0)}=\ket{0}\otimes\ket{0}$. (b) Same analysis as in (a) but considering as initial states the coherent states $\alpha_{1}(0)=\alpha_{1}$, $\alpha_{2}(0)=\alpha_{2}$ and |$\ket{\psi(0)}=\ket{\alpha_{1}}\otimes\ket{\alpha_{1}}$, where $\alpha_{1}=\langle\hat{a}_{1}\rangle$ and $\alpha_{2}=\langle\hat{a}_{2}\rangle$ are computed with the TWA in the steady state. (c) Same analysis as in (a) but for the photon number statistics $\delta n_{\ell}$. (d) Same analysis as in (b) but for the photon number statistics $\delta n_{\ell}$. ED and TWA show an excellent agreement in the vacuum, chaotic and high-photon number phases. The other parameters are set as in Fig. 2.
@@
In Fig. 10 we present the comparison between the TWA and the QCE. Figs. 10 (a) and (b) show $n_{L}$ and $\delta n_{L}$ computed solving cumulant equations truncated at order two. To compute the steady-state photon number and its fluctuations at each point in the phase diagram, we let the QCE dynamics evolve to the steady state, and we subsequently averaged the dynamics over a time window equal to $\Delta\tau=10^{3}$. The black line indicates the boundaries of the chaotic region computed from TWA equations. Data show a good qualitative agreement between TWA and QCE. However, as it is already evident from a rapid comparison between Figs. 2 (b) and (c) and Fig. 10, the two methods do not show a quantitative agreement.
@@
Figure 11: Saturation value of the steady-state phase OTOC, $D_{1,\ell}(\tau\to\infty)$ computed at different representative sites $\ell$ across the chain, as a function of the drive amplitude $F$ and the chain’s length $L$: (a) $\ell=1$, (b) $\ell=L/2$ , and (c) $\ell=L$ [same data as in Fig. 2 (c)]. The other parameters are set as in Fig. 2.
In this section, we provide additional plots to support the claim that the presence of semiclassical chaos is spread over the whole chain, and not only at the right boundary $\ell=L$ which is used in the main manuscript as a representative site. In Fig. 11, we plot the saturation value of the phase OTOC $D_{1,\ell}(\tau\to\infty)$ as a function of the drive amplitude $F$ and the chain’s length $L$, when evaluated at three representative sites across the chain: first $\ell=1$, middle $\ell=L/2$, and last site $\ell=L$ (that was already presented in the main text).
When $\ell=1$, the saturation value never reaches 1, signaling that the first site does not reach maximal decorrelation. This is in agreement with the OTOC’s dynamics studied in Fig. 8 (a) and (b) for a specific value of $F$ and $L$. The phase diagram for $\ell=L/2$ is instead essentially identical to the one realized for $\ell=L$. This analysis shows that the prethermal and thermal domains are equally chaotic.
@@
In Fig. 12, we compare the boundary-driven boundary-dissipative chain studied in this paper and the uniformly-driven uniformly-dissipative chain, studied, e.g., in Ref. [66], showing that semiclassical chaos is a feature peculiar of boundary drive and dissipation mechanisms. We focus on the last cavities of the two chains in the steady-state regime. In Fig. 12 (a), we show that the photon number $n_{L}$ exhibits a similar behavior with an abrupt jump around the same critical drive amplitude $F$. In Fig. 12 (b), we present the photon number statistics $\delta n_{L}$. In the case of the uniformly-driven uniformly-dissipative chain, large fluctuations are concentrated at the value of $F$ where the jump occurs, as expected. In contrast, in the case of boundary drive and dissipation, large fluctuations are spread over a much larger region. In Fig. 12 (c), we plot the saturation value of the steady-state semiclassical phase OTOC defined in Eq. (6), showing final evidence that only the boundary-driven, boundary-dissipative lattice hosts semiclassical chaos, while the dynamics in the uniformly-driven dissipative chain is regular.
Classical thermodynamics within the chaotic chain: application of the equipartition theorem
Figure 13: Comparison between the TWA equations (Truncated Wigner approximation) and the classical equations of motion (Gross-Pitaevksii classical equations of motion) in the chaotic regime. (a) TWA (blue line) and classical (black-dashed line) steady-state photon number $n_{\ell}$ as a function of $\ell$ for a $L=400$ chain. (b) Same as in panel (a) but for the circular variance $\Delta\varphi_{\ell}$. (c-e) Local momentum distribution $\textrm{P}(p_{\ell})$ computed from the TWA (blue lines) and the mean-field (gray line) equations. The site indices are (c) $\ell=5$, (d) $\ell=200$, (e) $\ell=390$. The dashed lines curves in panel (e) are the Maxwell-Boltzmann distributions associated with the TWA temperature (black-dashed line) and the classical temperature (red-dashed line). Results are computed by averaging over $N_{\rm traj}=10^{2}$ independent Wigner trajectories and over a time window $\Delta\tau=10^{4}$ once the steady state is reached. The drive amplitude is fixed to $F=7.5$. Results are computed by averaging over $N_{\rm traj}=10^{2}$ independent Wigner trajectories and over a time window $\Delta\tau=10^{4}$ once the steady state is reached. The other parameters are set as in Fig. 2.
Figure 14: Application of the equipartition theorem in the chaotic regime. (a) Temperature profile $T_{\ell}$ as a function of $\ell$ in a $L=400$ chain computed according to the equipartition theorem $T_{\ell}=|\Delta|\langle p_{\ell}^{2}\rangle$ in at low drive amplitude $F=5.5$ (pink line) and intermediate drive amplitude $F=7.5$ (purple line). (b-e) Momentum distributions $\textrm{P}(p_{\ell})$ as a function of $p_{\ell}$ (orange lines) compared to the thermal Maxwell-Boltzmann distribution in Eq. (40) with $T_{\ell}=|\Delta|\langle p_{\ell}^{2}\rangle$ at low drive amplitude $F=5.5$ at the sites (b) $\ell=50$, (c) $\ell=200$, (d) $\ell=340$, and (e) $\ell=370$. (f-i) Same as in panels (b-e) but for intermediate drive amplitude $F=7.5$. The other parameters are set as in Fig. 2.
The goal of this section is to discuss the similarities and differences between the TWA and classical Gross-Pitaevskii approaches in the chaotic regime, and the applicability of the equipartition theorem, a result derived from classical thermodynamics and largely applied in Ref. [50]. The equipartition theorem states that at thermal equilibrium, each quadratic degree of freedom contributes $T/2$ ($k_{B}=1$) to the average energy of the system. The equipartition theorem holds if: (i) quantum effects are negligible and the system is well described by classical mechanics, (ii) the energy is quadratic in the generalized coordinates like momentum and position, (iii) the system reached thermal equilibrium. For the Bose-Hubbard model, equipartition theorem does not apply in principle, even in semiclassical regimes, because of the presence of the Kerr term. In very diluite limits, however, nonlinearities play a negligible role and only quadratic terms contribute to the average energy. In Eq. (2), $\Delta$ fixes the energy scale of the quadratic Hamiltonian, and assuming that the Kerr term is negligible, one obtains that the distribution of momenta $\hat{p}:=i(\hat{a}^{\dagger}-\hat{a})/\sqrt{2}$ follows the thermal Maxwell-Boltzmann distribution
$\textrm{P}_{\rm MB}(p)=\sqrt{\frac{|\Delta|}{2\pi T}}e^{-|\Delta|p^{2}/2T}.$ (40)
The temperature can be estimated from the second moment of the distribution as $T=|\Delta|\langle p^{2}\rangle$. Within the TWA framework, $\langle p\rangle$ and $\langle p^{2}\rangle$ are given by $\sqrt{2}\langle\textrm{Im}(\alpha)\rangle$ and $2\langle\textrm{Im}^{2}(\alpha)\rangle$, respectively.
We start by studying the comparison between the TWA and Gross-Pitaevskii solutions for the chaotic regime. Results are collected in Fig. 13 for a $L=400$ chain at intermediate drive amplitude, $F=7.5$. In Figs. 13 (a) and (b) we present the photon number $n_{\ell}$ and the circular variance $\Delta\varphi_{\ell}$ as a function of $\ell$. Both quantities show the qualitative agreement between the two approaches: the circular variance rapidly saturates to one and the photon number remains flat in the bulk of the chain. The three distinct domains, nonthermal, prethermal and thermal are captured by both the approaches, as indicated by the momentum distributions in Figs. 13 (c-e). This indicates that in the chaotic regime, quantum fluctuations have a marginal role and the relevant physical phenomena can be attributed to classical nonlinear fluctuations.
We then apply the equipartition theorem. In Fig. 14 (a) we show the temperature estimated with the equipartition theorem in the chaotic regime for $F=5.5$ (pink line) and $F=7.5$ (purple line) in a $L=400$ chain. At $F=5.5$, $T_{\ell}$ decreases monotonically throughout the chain, in agreement with the behavior of the entropy density in Fig. 4 (c). Moreover, the momentum distributions $\textrm{P}(p_{\ell})$ reported in Figs. 14 (b-e) for $\ell=50$, $\ell=200$, $\ell=340$ and $\ell=370$ show a Gaussian profile, in agreement with the Maxwell-Boltzmann distribution in Eq. (40). At low drive amplitude, the effect of the nonlinear term reduces to breaking the model’s integrability, without shaping the local states. At $F=7.5$, $T_{\ell}$ saturates in the bulk of the chain, and decreases at the end of the chain, in correspondence of the thermal domain. This behavior is in sharp contrast with the temperature profile estimated with the Gibbs state, reported in Fig. 4 (a), and with the non-monotonic behavior of the steady-state entropy in Fig. 4 (c). As expected, the momentum distributions $\textrm{P}(p_{\ell})$ reported in Figs. 14 (f-i) for $\ell=50$, $\ell=200$, $\ell=340$ and $\ell=370$ show a strong non-Gaussian behavior in the bulk of the chain. At intermediate drive amplitudes, nonlinear fluctuations not only induce chaotic behavior, but also shape significantly the local states throughout the chain. The equipartition theorem breaks down because of the non negligible Kerr nonlinearity, activated by the stronger driving field. These nonlinear contributions are instead captured by the Gibbs ansatz, which unveils the anomalous heating in the prethermal domain. In the thermal domain ($\ell=370$), the shape of $\textrm{P}(p_{\ell})$ has a bell shape and its profile is reproduced by a thermal Maxwell-Boltzmann distribution, with a small deviation due again to the Kerr nonlinearity.
@@
Modeling the local state of the chaotic Bose-Hubbard chain with $\mathbb{U}(1)$-symmetric impurity models
@@
We consider two single-site driven-dissipative impurity models that describe $\mathbb{U}(1)$-symmetric quantum states exhibiting, across their parameter space, thermal features and population inversion. Population inversion is achieved when higher energy states are more populated than lower energy states, in contrast to a thermal population, and requires an external energy source [100, 85, 101]. Population inversion is an essential ingredient to obtain lasing states, where stimulated emission dominates over absorption [102]. The Hamiltonian of both the models is taken as the local Hamiltonian of the Bose-Hubbard chain in Eq. (2), $\hat{H}_{\rm imp}=\omega_{0}\hat{a}^{\dagger}\hat{a}+U\hat{a}^{\dagger 2}\hat{a}^{2}/2$. The first impurity model, which we refer as the 2-photon decay model, is defined by the set of Lindblad dissipators described in the main text,
$\hat{L}_{\uparrow}=\sqrt{\gamma^{\uparrow}}\hat{a}^{\dagger},\qquad\hat{L}_{\downarrow}=\sqrt{\gamma^{\downarrow}}\hat{a},\qquad\hat{L}_{\phi}=\sqrt{\gamma^{\phi}}\hat{a}\hat{a}^{\dagger},\qquad\hat{L}_{\rm s}=\sqrt{\gamma^{\rm s}}\hat{a}^{2}\,.$ (41)